OpenStudy (zzr0ck3r):

Fun set theory question for budding set theorist. Let \(B^A\) be the set of all functions from \(A\) to \(B:=\{0,1\}\). Show that \(\mathcal{P}(A)\) has the same cardinality as \(B^A\) for any set \(A\).

OpenStudy (anonymous):

Well, we know in the finite case \(|\mathcal P (A)|=2^{|A|}\) and \(|B^A| = |B|^{|A|} = 2^{|A|}\)

OpenStudy (anonymous):

I think this applies when \(|A|>\aleph_i\).

OpenStudy (zzr0ck3r):

yeah it does.

OpenStudy (zzr0ck3r):

A can have any cardinality

OpenStudy (anonymous):

You could create a homomorphism as well. \[ \forall S\subseteq A: \exists f\in B^A: f(x)= \begin{cases}1 &x\in S \\ 0 &x\notin S\end{cases} \]

OpenStudy (anonymous):

You can build a unique \(f\) from \(S\) and vise versa.

OpenStudy (anonymous):

yeah, the bijection is easy -- for \(f\in B^A\) we associate the inverse image \(f^{-1}(1)\in 2^A\), and for \(S\in2^A\) we associate \(f\) such that \(f[S]=1,f[S^c]=0\)

OpenStudy (anonymous):

this is a side-effect of the fact that sets of things in \(A\) are wholly characterized by their members (i.e. for what \(a\in A\) we have that \(a\in S\in 2^A\)), and membership is a binary map \(S\to \{0,1\}\)